# Bilevel ZOFO: Bridging Parameter-Efficient and Zeroth-Order Techniques for Efficient LLM Fine-Tuning and Meta-Training

### v2

@@ 1 Introduction
Fine-tuning pretrained Large Language Models (LLMs) has become a standard approach for downstream tasks. Traditional First-Order (FO) optimizers like Adam [20], commonly used for this process, rely on backpropagation. However, as highlighted in Malladi et al. [33], computing gradients for LLMs can require up to 12 times the memory needed for inference. This scaling challenge becomes even more pronounced as models grow larger, imposing significant memory demands and complicating the fine-tuning process, especially in resource-constrained environments.
To address these computational challenges, Parameter-Efficient Fine-Tuning (PEFT) methods have been developed. These techniques freeze most of the model’s parameters and train only a small subset, significantly reducing both memory and computational overhead. Popular PEFT approaches include prompt tuning, LoRA fine-tuning, and prefix tuning. Prompt tuning [22, 38, 56, 46] optimizes continuous prompt vectors that are concatenated with the input embeddings, while prefix tuning [24] introduces learnable prefix tokens that serve as conditioning variables at each transformer layer. LoRA (Low-Rank Adaptation) [16, 15] modifies the model’s attention and feedforward layers by injecting low-rank trainable matrices, further reducing the resources required for fine-tuning.
@@
While Parameter-Efficient Fine-Tuning (PEFT) methods reduce training costs and memory usage, they may not always achieve the same level of task-specific performance as full model fine-tuning. Research has shown that for tasks requiring high accuracy, complex adaptations, or domain-specific knowledge, full fine-tuning often outperforms PEFT approaches due to its ability to adjust all model parameters for better adaptation [16, 24, 57]. To make full model fine-tuning more computationally feasible, Zeroth-Order (ZO) methods offer an alternative by reducing the high computational cost. Rather than computing gradients via backpropagation, zeroth-order methods estimate the gradient using only the forward pass. Initially explored in the 1990s [48, 36, 11, 6, 28], these methods have recently gained traction for fine-tuning LLMs [33, 4, 25] and have been shown to be able to outperform FO PEFT methods given enough training time [61].
A major limitation of ZO methods is their slow convergence due to the need for gradient estimation. For instance, MeZO [33] required 10 times more iterations than PEFT baselines to match or exceed their performance. Additionally, ZO methods suffer from extreme sensitivity to prompt selection. In tasks like sentiment analysis with the SST-2 dataset, templated prompts (e.g., <CLS><CLS> text data. It was [terrible $\mid$ great]. <SEP><SEP>") are crucial for success [61]. These prompts effectively align the text data with task-specific objectives. As a result, prompt selection becomes an important hyperparameter that can significantly affect performance. In particular, ZO methods have been shown to be highly sensitive to prompt selections  [33]. Without proper prompts, the performance of MeZO can drastically drop (Table 1).
@@ 2.1 Zeroth-Order Methods in Fine-Tuning LLMs
MeZO [33] pioneered zeroth-order (ZO) fine-tuning for LLMs, demonstrating compatibility with both full-model and PEFT approaches while improving computational efficiency. Subsequent work provided a benchmark for ZO optimization methods [61], and expanded ZO applications to variance reduction [10], federated fine-tuning [39, 25], softmax layers [5], sparse tuning [14, 29], and privacy [49]. In contrast, we propose a bilevel training framework that unifies the strengths of ZO full-model and FO PEFT fine-tuning, outperforming both individually, while using as much resources.
@@ 2.2 First-Order Methods for Bilevel Optimization
Bilevel optimization is computationally demanding, especially for LLMs, due to the cost of computing hypergradients. Classical approaches rely on second-order methods [8, 9, 7, 23, 41, 12, 2, 30], while recent work [31, 44, 27, 21, 26, 18, 32] bypass the need for second-order information by reformulating the bilevel problem as a constrained optimization problem. We build on this by incorporating ZO approximations into the upper-level optimization to bypass full gradient computation for LLMs. These methods significantly reduce computational costs by eliminating the need for second-order information. Nevertheless, when fine tuning LLMs, back propagation for calculating the gradient of an LLM is still too expensive. Liu et al. [27] and Lu and Mei [31] explore the convergence of their proposed methods to the original bilevel problem, while other approaches only demonstrate convergence to the penalized problem. In this paper, we adapt the method from Lu and Mei [31] to approximate part of the upper-level parameters using a ZO approximation in order to address the challenge posed by the large number of training parameters in large language models. We also provide convergence guarantees for this adapted zeroth-order-first-order method.
@@ 3.1 Preliminaries and Notation
Where $\mathbf{p}$} corresponds to the embeddings of the hard prompt (as shown in Table 13 in the appendix of [33]), the model above reduces to classical fine-tuning on a single downstream task. In model (1), the parameters of the PEFT model, $\mathbf{p}$}, are fixed.
@@ 3.2 Bilevel Model
To reduce the computational cost, following [31], we consider using a penalty method for the bilevel problem (2). Specifically, (2) is equivalent to the following constrained optimization problem:
@@
requires calculating the gradient with respect to $\bm{\theta}$}, i.e, $\nabla_{{\bm{\theta}}}F({\bm{\theta}}^{k},{\mathbf{p}}^{k};\mathfrak{D}_{f})$. Given the large scale of $\bm{\theta}$} in LLMs, this is computationally expensive. To avoid this, we use zeroth-order (ZO) information to approximate the gradient $\nabla_{\bm{\theta}}G$. Following [33, 61, 14], we employ the Simultaneous Perturbation Stochastic Approximation (SPSA) as a classical zeroth-order gradient estimator. Specifically, at each iteration $k$, we sample ${\mathbf{z}}^{k}\sim N(0,I_{d})$, recalling that $d$ is the dimension of $\bm{\theta}$}. We then approximate the gradient $\nabla_{\bm{\theta}}F$ as follows:
@@
This approach leads to the final algorithm (Algorithm 2 and Figure 5) for fine-tuning LLMs using the bilevel model (2). We refer to this method as the Bilevel Zeroth-Order-First-Order (Bilevel ZOFO) method. In Appendix B, we show that Bilevel ZOFO converges at a rate of $O(\epsilon^{-2})$ under widely accepted assumptions. [^2] In our theorem, we assume strong convexity on the lower level objective function, as is common in other theoretical work on bilevel optimization. While strong convexity facilitates theoretical analysis, our experiments demonstrate that the method remains robust even when this condition is not strictly satisfied. The complexity of Bilevel ZOFO matches that in previous ZO minimax algorithm in [53] but solves our bilevel optimization problem (2) and does not depend on the dimensionality $d$ thanks to the efficient rank assumption B.5, providing efficiency guarantee for our algorithm.
@@ 4 Experiments
We conduct extensive experiments on various LLMs of different scales to demonstrate the effectiveness of bilevel-ZOFO in improving current zeroth order methods and PEFT. We also conduct experiments in testing its potential in meta training. Noticing that our proposed structure is able to incorporate any variation of zeroth-order methods in the upper-level step and any PEFT method in the lower level, to maintain focus on testing the effectiveness of the proposed bilevel structure and its unique multitask learning capabilities, we used the classic MeZO  [33].
@@ 4.1.1 Experimental Setting
Following MeZO [33], we evaluate our approach on a range of classification and multiple-choice tasks. In this setting, training and testing are conducted on the same task. We employ prompt-tuning [22], prefix-tuning [24], and LoRA [16]- well-known PEFT baselines-for lower-level training to validate bilevel-ZOFO under different conditions and resource constraints. During each lower-level update, we update only the PEFT parameters, and during the upper-level optimization step, we tune the full model using zeroth-order gradient approximation. We perform 10 lower-level updates between each pair of upper-level updates. For each task, we randomly sample 1000 examples for training, 500 examples for validation, and 1000 examples for testing. We use the Adam optimizer [20] and report test accuracy or F1-score.
@@ 4.1.2 Results
Following the setup of Table 5 in the MeZO paper [33], we evaluate both MeZO and Bilevel-ZOFO in two scenarios: 1- where a simple task prompt is prepended to each input versus 2- where no such prompt is used. Table 1 reports results for tuning OPT-1.3B on SST-2 and COPA using LoRA as the PEFT method. Our findings show that Bilevel-ZOFO is markedly less sensitive to prompt variations than MeZO; the performance gap between prompted and unprompted settings is substantially smaller for Bilevel-ZOFO.
@@
Table 2 presents the test metrics when applying bilevel-ZOFO and baselines to fine-tune OPT-1.3B [60] on a downstream task. Table 3 demonstrates the results for Llama2-7b [50]. We can make the following observations:
@@ 4.1.3 Memory Profiling and Wall Clock Time Analysis
We also present a wall-clock time analysis of bilevel-ZOFO compared to the baseline. As shown in Table 4, similar to MeZO [33], we observe that zeroth-order steps exhibit higher latency compared to first-order steps. The results indicate that our bilevel-ZOFO achieves comparable delays to the FO-PEFT method while significantly reducing step duration compared to MeZO. Moreover, as highlighted in Table 2, bilevel-ZOFO outperforms both methods on average.
@@ 4.3 Adopting Bilevel ZOFO to Meta-learning
Following Min et al. [34], we evaluate the performance of bilevel-ZOFO as a fast and efficient meta-learning algorithm. We perform experiments using four of the distinct meta-learning settings: classification-to-classification, non-classification-to-classification, QA-to-QA, and non-QA-to-QA. For instance, in non-classification-to-classification setting, we train on a number of non-classification subtasks and test on a number of distinct classification subtasks. Each of these meta-learning tasks includes a set of training sub-tasks and a different set of test sub-tasks. The sub-tasks are sourced from CROSSFIT [55] and UNIFIEDQA [19], comprising a total of 142 unique sub-tasks. These sub-tasks cover a variety of problems, including text classification and question answering all in English. We use GPT2-Large [40] as the base model for these experiments.
@@
  • MetaICL [34]: A method for meta-learning with in-context learning. MetaICL tunes all the parameters of the base model using the first-order method. In both training and testing, the model is given $k$ demonstration examples, $(a_{1},b_{1}),\dots,(a_{k},b_{k})$}, where $b_{i}$ represents either classification labels or possible answers in question-answering tasks, along with one test example $(a,b)$. The input is formed by concatenating the demonstration examples $a_{1},b_{1},\dots,a_{k},b_{k},a$. The model then computes the conditional probability of each label, and the label with the highest probability is selected as the prediction.
@@
We sample 768 examples from each training sub-task. We train MetaICL in their original setting for 30,000 steps. To train our method, we split the training dataset of each sub-task to two subsets, 256 samples as the development dataset for upper-level updates and 512 samples for lower-level training. For each outer iteration of our method, we randomly sample a subset of 5 training tasks. We perform 10 lower-level updates between each pair of upper-level updates. To keep bilevel-ZOFO as lightweight as possible, unlike MetaICL, we DO NOT include demonstration examples in the inputs. Since bilevel-ZOFO uses significantly less memory and has much faster updates compared to MetaICL, theoretically we are able to train it for many more iterations within the same total training duration as MetaICL. However, due to resource constraints, we only train bilevel-ZOFO for 50,000 iterations. Similar to  [33], we did not observe a plateau in performance for bilevel-ZOFO, indicating that further training can yield additional improvements.
For both ICL and MetaICL, during the testing phase the model is given $k=4$ demonstration examples for each test data point. We don’t use demonstration examples in test samples for bilevel-ZOFO evaluation. We evaluate the zero-shot capabilities of our method as well as the performance of the final model LoRA-tuned for 10 additional iterations on 4 demonstration samples from each class of each test sub-task. Similar to [34], we report Macro-averaged F1 as the evaluation metric. See Appendix D.3 for all training details.
@@ A.1 Zeroth order in fine tuning LLMs
MeZO [33] is the first work to use Zeroth-Order (ZO) methods to finetune LLMs for downstream tasks. They demonstrate that their method is compatible with both full-parameter tuning and parameter-efficient tuning techniques, such as LoRA and prefix tuning, while being significantly more computationally efficient. Zhang et al. [61] provide a benchmark for ZO optimization in the context of LLM fine-tuning, comparing different ZO optimizers and applying the method to various models. Gautam et al. [10] introduce variance reduction techniques into ZO methods for fine-tuning, improving both stability and convergence. In addition, ZO methods are applied in federated fine-tuning by Qin et al. [39] and Ling et al. [25]. Deng et al. [5] implement ZO optimization for softmax units in LLMs. Guo et al. [14] and Liu et al. [29] explore fine-tuning a minimal subset of LLM parameters using ZO methods by sparsifying gradient approximation or the perturbations used in gradient estimation. Tang et al. [49] investigate the privacy of ZO optimization methods.

### v3

# Bilevel ZOFO: Efficient LLM Fine-Tuning and Meta-Training

@@ 1 Introduction
Fine-tuning pretrained Large Language Models (LLMs) has become a standard approach for downstream tasks. Traditional First-Order (FO) optimizers like Adam [AdamKingmaB14], commonly used for this process, rely on backpropagation. However, as highlighted in MalladiGNDL0A23Mezo, computing gradients for LLMs can require up to 12 times the memory needed for inference. This scaling challenge becomes even more pronounced as models grow larger, imposing significant memory demands and complicating the fine-tuning process, especially in resource-constrained environments.
To address these computational challenges, Parameter-Efficient Fine-Tuning (PEFT) methods have been developed. These techniques freeze most of the model’s parameters and train only a small subset, significantly reducing both memory and computational overhead. Popular PEFT approaches include prompt tuning, LoRA fine-tuning, and prefix tuning. Prompt tuning [LesterAC21PromptTuning, QinE21PromptTuning, YuCL023, shin-prompt-sensitivity] optimizes continuous prompt vectors that are concatenated with the input embeddings, while prefix tuning [LiL20PrefixTuning] introduces learnable prefix tokens that serve as conditioning variables at each transformer layer. LoRA (Low-Rank Adaptation) [HuSWALWWC22LORA, HoulsbyGJMLGAG19LORA] modifies the model’s attention and feedforward layers by injecting low-rank trainable matrices, further reducing the resources required for fine-tuning.
@@
While Parameter-Efficient Fine-Tuning (PEFT) methods reduce training costs and memory usage, they may not always achieve the same level of task-specific performance as full model fine-tuning. Research has shown that for tasks requiring high accuracy, complex adaptations, or domain-specific knowledge, full fine-tuning often outperforms PEFT approaches due to its ability to adjust all model parameters for better adaptation [HuSWALWWC22LORA, LiL20PrefixTuning, ZakenGR22]. To make full model fine-tuning more computationally feasible, Zeroth-Order (ZO) methods offer an alternative by reducing the high computational cost. Rather than computing gradients via backpropagation, zeroth-order methods estimate the gradient using only the forward pass. Initially explored in the 1990s [Spall92, NesterovS17, GhadimiL13a, DuchiJWW15, LiuCKZHV20], these methods have recently gained traction for fine-tuning LLMs [MalladiGNDL0A23Mezo, DeLiMaSo23, LingCYL024] and have been shown to be able to outperform FO PEFT methods given enough training time [ZhangLHLZZCLY0W24Zobench].
A major limitation of ZO methods is their slow convergence due to the need for gradient estimation. For instance, MeZO [MalladiGNDL0A23Mezo] required 10 times more iterations than PEFT baselines to match or exceed their performance. Additionally, ZO methods suffer from extreme sensitivity to prompt selection. In tasks like sentiment analysis with the SST-2 dataset, templated prompts (e.g., <CLS><CLS> text data. It was [terrible $\mid$ great]. <SEP><SEP>") are crucial for success [ZhangLHLZZCLY0W24Zobench]. These prompts effectively align the text data with task-specific objectives. As a result, prompt selection becomes an important hyperparameter that can significantly affect performance. In particular, ZO methods have been shown to be highly sensitive to prompt selections  [MalladiGNDL0A23Mezo]. Without proper prompts, the performance of MeZO can drastically drop (Table 1).
@@ 2.1 Zeroth-Order Methods in Fine-Tuning LLMs
MeZO [MalladiGNDL0A23Mezo] pioneered zeroth-order (ZO) fine-tuning for LLMs, demonstrating compatibility with both full-model and PEFT approaches while improving computational efficiency. Subsequent work provided a benchmark for ZO optimization methods [ZhangLHLZZCLY0W24Zobench], and expanded ZO applications to variance reduction [GautamPZRH24], federated fine-tuning [QinCQDLD24, LingCYL024], softmax layers [DengLMS24], sparse tuning [GuoLZLY24, LiuZGCHY24], and privacy [TaPNMM24]. In contrast, we propose a bilevel training framework that unifies the strengths of ZO full-model and FO PEFT fine-tuning, outperforming both individually, while using as much resources.
@@ 2.2 First-Order Methods for Bilevel Optimization
Bilevel optimization is computationally demanding, especially for LLMs, due to the cost of computing hypergradients. Classical approaches rely on second-order methods [FrDoFrPo17, FrFrSaGrPo18, FiAbLe17, LiGuHu21, RaFiKaLe19, GhWa18, ChKaZh22, LiXiFeZhYoPiUZ18], while recent work [LuMei24, ShenC23, LiuLYZZ24, KwonKWN23, LiuYWSL22, ji-bilevel-rebuttal-1, mackay-bilevel-rebuttal-2] bypass the need for second-order information by reformulating the bilevel problem as a constrained optimization problem. We build on this by incorporating ZO approximations into the upper-level optimization to bypass full gradient computation for LLMs. These methods significantly reduce computational costs by eliminating the need for second-order information. Nevertheless, when fine tuning LLMs, back propagation for calculating the gradient of an LLM is still too expensive. LiuLYZZ24 and LuMei24 explore the convergence of their proposed methods to the original bilevel problem, while other approaches only demonstrate convergence to the penalized problem. In this paper, we adapt the method from LuMei24 to approximate part of the upper-level parameters using a ZO approximation in order to address the challenge posed by the large number of training parameters in large language models. We also provide convergence guarantees for this adapted zeroth-order-first-order method.
@@ 3.1 Preliminaries and Notation
Where $\mathbf{p}$} corresponds to the embeddings of the hard prompt (as shown in Table 13 in the appendix of [MalladiGNDL0A23Mezo]), the model above reduces to classical fine-tuning on a single downstream task. In model (1), the parameters of the PEFT model, $\mathbf{p}$}, are fixed.
@@ 3.2 Bilevel Model
To reduce the computational cost, following [LuMei24], we consider using a penalty method for the bilevel problem (2). Specifically, (2) is equivalent to the following constrained optimization problem:
@@
requires calculating the gradient with respect to $\bm{\theta}$}, i.e, $\nabla_{{\bm{\theta}}}F({\bm{\theta}}^{k},{\mathbf{p}}^{k};\mathfrak{D}_{f})$. Given the large scale of $\bm{\theta}$} in LLMs, this is computationally expensive. To avoid this, we use zeroth-order (ZO) information to approximate the gradient $\nabla_{\bm{\theta}}G$. Following [MalladiGNDL0A23Mezo, ZhangLHLZZCLY0W24Zobench, GuoLZLY24], we employ the Simultaneous Perturbation Stochastic Approximation (SPSA) as a classical zeroth-order gradient estimator. Specifically, at each iteration $k$, we sample ${\mathbf{z}}^{k}\sim N(0,I_{d})$, recalling that $d$ is the dimension of $\bm{\theta}$}. We then approximate the gradient $\nabla_{\bm{\theta}}F$ as follows:
@@
This approach leads to the final algorithm (Algorithm 2 and Figure 5) for fine-tuning LLMs using the bilevel model (2). We refer to this method as the Bilevel Zeroth-Order-First-Order (Bilevel ZOFO) method. In Appendix B, we show that Bilevel ZOFO converges at a rate of $O(\epsilon^{-2})$ under widely accepted assumptions. [^2] In our theorem, we assume strong convexity on the lower level objective function, as is common in other theoretical work on bilevel optimization. While strong convexity facilitates theoretical analysis, our experiments demonstrate that the method remains robust even when this condition is not strictly satisfied. The complexity of Bilevel ZOFO matches that in previous ZO minimax algorithm in [wang2023zeroth] but solves our bilevel optimization problem (2) and does not depend on the dimensionality $d$ thanks to the efficient rank assumption B.5, providing efficiency guarantee for our algorithm.
@@ 4 Experiments
We conduct extensive experiments on various LLMs of different scales to demonstrate the effectiveness of bilevel-ZOFO in improving current zeroth order methods and PEFT. We also conduct experiments in testing its potential in meta training. Noticing that our proposed structure is able to incorporate any variation of zeroth-order methods in the upper-level step and any PEFT method in the lower level, to maintain focus on testing the effectiveness of the proposed bilevel structure and its unique multitask learning capabilities, we used the classic MeZO  [MalladiGNDL0A23Mezo].
@@ 4.1.1 Experimental Setting
Following MeZO [MalladiGNDL0A23Mezo], we evaluate our approach on a range of classification and multiple-choice tasks. In this setting, training and testing are conducted on the same task. We employ prompt-tuning [LesterAC21PromptTuning], prefix-tuning [LiL20PrefixTuning], and LoRA [HuSWALWWC22LORA]- well-known PEFT baselines-for lower-level training to validate bilevel-ZOFO under different conditions and resource constraints. During each lower-level update, we update only the PEFT parameters, and during the upper-level optimization step, we tune the full model using zeroth-order gradient approximation. We perform 10 lower-level updates between each pair of upper-level updates. For each task, we randomly sample 1000 examples for training, 500 examples for validation, and 1000 examples for testing. We use the Adam optimizer [AdamKingmaB14] and report test accuracy or F1-score.
@@ 4.1.2 Results
Following the setup of Table 5 in the MeZO paper [MalladiGNDL0A23Mezo], we evaluate both MeZO and Bilevel-ZOFO in two scenarios: 1- where a simple task prompt is prepended to each input versus 2- where no such prompt is used. Table 1 reports results for tuning OPT-1.3B on SST-2 and COPA using LoRA as the PEFT method. Our findings show that Bilevel-ZOFO is markedly less sensitive to prompt variations than MeZO; the performance gap between prompted and unprompted settings is substantially smaller for Bilevel-ZOFO.
@@
Table 2 presents the test metrics when applying bilevel-ZOFO and baselines to fine-tune OPT-1.3B [OPT] on a downstream task. Table 3 demonstrates the results for Llama2-7b [Llama2]. We can make the following observations:
@@ 4.1.3 Memory Profiling and Wall Clock Time Analysis
We also present a wall-clock time analysis of bilevel-ZOFO compared to the baseline. As shown in Table 4, similar to MeZO [MalladiGNDL0A23Mezo], we observe that zeroth-order steps exhibit higher latency compared to first-order steps. The results indicate that our bilevel-ZOFO achieves comparable delays to the FO-PEFT method while significantly reducing step duration compared to MeZO. Moreover, as highlighted in Table 2, bilevel-ZOFO outperforms both methods on average.
@@ 4.3 Adopting Bilevel ZOFO to Meta-learning
Following MinLZH22MetaICL, we evaluate the performance of bilevel-ZOFO as a fast and efficient meta-learning algorithm. We perform experiments using four of the distinct meta-learning settings: classification-to-classification, non-classification-to-classification, QA-to-QA, and non-QA-to-QA. For instance, in non-classification-to-classification setting, we train on a number of non-classification subtasks and test on a number of distinct classification subtasks. Each of these meta-learning tasks includes a set of training sub-tasks and a different set of test sub-tasks. The sub-tasks are sourced from CROSSFIT [YeLR21CrossfitFewShot] and UNIFIEDQA [KhashabiMKSTCH20], comprising a total of 142 unique sub-tasks. These sub-tasks cover a variety of problems, including text classification and question answering all in English. We use GPT2-Large [radford2019gpt2] as the base model for these experiments.
@@
  • MetaICL [MinLZH22MetaICL]: A method for meta-learning with in-context learning. MetaICL tunes all the parameters of the base model using the first-order method. In both training and testing, the model is given $k$ demonstration examples, $(a_{1},b_{1}),\dots,(a_{k},b_{k})$}, where $b_{i}$ represents either classification labels or possible answers in question-answering tasks, along with one test example $(a,b)$. The input is formed by concatenating the demonstration examples $a_{1},b_{1},\dots,a_{k},b_{k},a$. The model then computes the conditional probability of each label, and the label with the highest probability is selected as the prediction.
@@
We sample 768 examples from each training sub-task. We train MetaICL in their original setting for 30,000 steps. To train our method, we split the training dataset of each sub-task to two subsets, 256 samples as the development dataset for upper-level updates and 512 samples for lower-level training. For each outer iteration of our method, we randomly sample a subset of 5 training tasks. We perform 10 lower-level updates between each pair of upper-level updates. To keep bilevel-ZOFO as lightweight as possible, unlike MetaICL, we DO NOT include demonstration examples in the inputs. Since bilevel-ZOFO uses significantly less memory and has much faster updates compared to MetaICL, theoretically we are able to train it for many more iterations within the same total training duration as MetaICL. However, due to resource constraints, we only train bilevel-ZOFO for 50,000 iterations. Similar to  [MalladiGNDL0A23Mezo], we did not observe a plateau in performance for bilevel-ZOFO, indicating that further training can yield additional improvements.
For both ICL and MetaICL, during the testing phase the model is given $k=4$ demonstration examples for each test data point. We don’t use demonstration examples in test samples for bilevel-ZOFO evaluation. We evaluate the zero-shot capabilities of our method as well as the performance of the final model LoRA-tuned for 10 additional iterations on 4 demonstration samples from each class of each test sub-task. Similar to [MinLZH22MetaICL], we report Macro-averaged F1 as the evaluation metric. See Appendix D.3 for all training details.
@@ 6 Acknowledgments
Limitations and Broader Impact
This paper introduces an optimization framework that enhances the efficiency of fine-tuning large language models. By reducing computational costs and memory requirements, the approach promotes environmental sustainability and broadens access to advanced AI tools, promoting accessibility in AI development. While our framework is designed for scalability, we have not tested very large LLMs due to resource constraints. However, our experiments sufficiently validate the research idea. Future work includes exploring masked ZO tuning for efficiency and applying our approach to style mixing in image generation models.
@@ A.1 Zeroth order in fine tuning LLMs
MeZO [MalladiGNDL0A23Mezo] is the first work to use Zeroth-Order (ZO) methods to finetune LLMs for downstream tasks. They demonstrate that their method is compatible with both full-parameter tuning and parameter-efficient tuning techniques, such as LoRA and prefix tuning, while being significantly more computationally efficient. ZhangLHLZZCLY0W24Zobench provide a benchmark for ZO optimization in the context of LLM fine-tuning, comparing different ZO optimizers and applying the method to various models. GautamPZRH24 introduce variance reduction techniques into ZO methods for fine-tuning, improving both stability and convergence. In addition, ZO methods are applied in federated fine-tuning by QinCQDLD24 and LingCYL024. DengLMS24 implement ZO optimization for softmax units in LLMs. GuoLZLY24 and LiuZGCHY24 explore fine-tuning a minimal subset of LLM parameters using ZO methods by sparsifying gradient approximation or the perturbations used in gradient estimation. TaPNMM24 investigate the privacy of ZO optimization methods.